If the demand function of a commodity is Q = 36 − 4P, where P and Q are price and quantity respectively, determine the price elasticity of demand when the price is R5. Indicate whether demand is elastic or inelastic at this price and provide justification for your answer. [1] εd = −1,25; because | − 1,25| = 1,25 > 1, demand is elastic [2] εd = −0,25; because | − 0,25| = −0,25 < 1, demand is elastic [3] εd = 0,25; because |0,25| = 0,25 < 1, demand is inelastic [4] εd = 1,25; because |1,25| = 1,25 > 1, demand is inelastic Explain


1
Expert's answer
2021-05-25T16:58:07-0400

Demand Function ="Q = 36-4P"

Price = 5

"Q = 36 - 4 \times 5"

  "= 36 - 20"

"=16"

"\frac{∆Q}{∆P}= -4"

Price Elasticity Of demand"=\frac{ P}{Q}\times \frac{\delta Q}{\delta P}"

"=\frac{5}{16}x-4"

"=-1.25"

Hence the option A is correct

"εd = −1,25; because | − 1,25| = 1,25 > 1, demand is elastic"